OpenStudy (anonymous):

Let f(x) = 12 over the quantity of 4 x + 2. Find f(-1). -6 -2 2 6 Can you plz help me out with this?@phi

OpenStudy (anonymous):

can u plz help? @IMStuck

OpenStudy (carolina4567):

Hey @IMStuck did u recieve my message cause openstudys not working lately ? But did u recieve it ?

OpenStudy (anonymous):

\[f(-1) = \frac{ 12 }{ 4 \times (-1) + 2 } = \frac {12} {-2} = -6\] The answer is -6

OpenStudy (anonymous):

thank u so much!

OpenStudy (imstuck):

I thought you were doing inverses! No wonder I couldn't get the answer! You also denote the inverse of a function with f(-1), but usually the (-1) is in the exponent position. This is much easier than I was making it!

OpenStudy (anonymous):

lol thats what i tought at first too